# Imperative Learning: A Self-supervised Neuro-Symbolic Learning Framework for Robot Autonomy

### v6

@@ 4 Applications and Examples
Additionally, since IL is a self-supervised learning framework consisting of three primary components, we have three kinds of learning types. This includes (A) given known (pre-trained or human-defined) reasoning engines such as logical reasoning, physical principles, and geometrical analytics, robots can learn a logic-, physics-, or geometry-aware neural perception system, respectively, in a self-supervised manner; (B) given neural perception systems such as a vision foundation model (Kirillov et al., 2023), robots can discover the world rules, e.g., traffic rules, and then apply the rules to future events; and (C) given a memory system, (e.g., experience, world rules, or maps), robots can simultaneously update the neural system and reasoning engine so that they can adapt to novel environments with a new set of rules. Note that if one of the modules has a good initialization, all the three modules can be further finetuned simultaneously. The five examples will cover all three learning types.
@@ 4.1 Closed-form Solution
where the LL solutions ${\bm{\mu}}^{*}({\bm{\theta}},{\bm{\gamma}})$ and ${\bm{\nu}}^{*}({\bm{\theta}},{\bm{\gamma}})$, that contain the implicit components $\frac{\partial{\bm{\mu}}^{*}}{\partial{\bm{\theta}}}$, $\frac{\partial{\bm{\mu}}^{*}}{\partial{\bm{\gamma}}}$, $\frac{\partial{\bm{\nu}}^{*}}{\partial{\bm{\theta}}}$, and $\frac{\partial{\bm{\nu}}^{*}}{\partial{\bm{\gamma}}}$, can be calculated directly due to the closed-form of ${\bm{\mu}}^{*}$ and ${\bm{\nu}}^{*}$. As a result, standard gradient-based algorithms can be applied in updating ${\bm{\theta}}$ and ${\bm{\gamma}}$ with gradients given by (2). In this case, there is no approximation error induced by the LL minimization, in contrast to existing widely-used implicit and unrolled differentiation methods that require ensuring a sufficiently small or decreasing LL optimization error for guaranteeing the convergence (Ji et al., 2021).
@@
Many symbolic costs can be effectively addressed through closed-form solutions. For example, both linear quadratic regulator (LQR) (Shaiju and Petersen, 2008) and Dijkstra’s algorithm (Dijkstra, 1959) can be solved with a determined optimal solution. To demonstrate the effectiveness of IL for closed-form solutions, we next present two examples in path planning to utilize the neural model for reducing the search and sampling space of symbolic optimization.
@@
The A algorithm is a graph search technique that aims to find the global shortest feasible path between two nodes (Hart et al., 1968). Specifically, A selects the path passing through the next node $n\in G$ that minimizes
@@
where $s(n)$ is the cost from the start node to $n$ and $h(n)$ is a heuristic function that predicts the cheapest path cost from $n$ to the goal. The heuristic cost can take the form of various metrics such as the Euclidean and Chebyshev distances. It can be proved that if the heuristic function $h(n)$ is admissible and monotone, i.e., $\forall m\in G,h(n)\leq h(m)+d(m,n)$, where $d(m,n)$ is the cost from $m$ to $n$, A is guaranteed to find the optimal path without searching any node more than once (Dechter and Pearl, 1985). Due to this optimality, A became one of the most widely used methods in path planning (Paden et al., 2016; Smith et al., 2012; Algfoor et al., 2015).
However, A encounters significant limitations, particularly in its requirement to explore a large number of potential nodes. This exhaustive search process can be excessively time-consuming, especially for low-power robot systems. To address this, recent advancements showed significant potential for enhancing efficiency by predicting a more accurate heuristic cost map using data-driven methods (Choudhury et al., 2018; Yonetani et al., 2021a; Kirilenko et al., 2023). Nevertheless, these algorithms utilize optimal paths as training labels, which face challenges in generalization, leading to a bias towards the patterns observed in the training datasets.
Figure 2: The framework of iA search. The network predicts a confined search space, leading to overall improved efficiency. The A search algorithm eliminates the label dependence, resulting in a self-supervised path planning framework.
@@
To address this limitation, we leverage the IL framework to remove dependence on path labeling. Specifically, we utilize a neural network $f$ to estimate the $h$ value of each node, which can be used to reduce the search space. Subsequently, we integrate a differentiable A module, serving as the symbolic reasoning engine $g$ in (1b), to determine the most efficient path. This results in an effective framework depicted in Figure 2, which we refer to as imperative A (iA) algorithm. Notably, the iA framework can operate on a self-supervised basis inherent in the IL framework, eliminating the need for annotated labels.
Specifically, the iA algorithm is to minimize
@@
As shown in (2), once the gradient calculation is completed, we backpropagate the cost $U$ directly to the network $f$. This process is facilitated by the closed-form solution of the A algorithm. For the sake of simplicity, we employ the differentiable A algorithm as introduced by Yonetani et al. (2021a). It transforms the node selection into an argsoftmax operation and reinterprets node expansion as a series of convolutions, leveraging efficient implementations and AutoDiff in PyTorch (Paszke et al., 2019).
@@
End-to-end local planning, which integrates perception and planning within a single model, has recently attracted considerable interest, particularly for its potential to enable efficient inference through data-driven methods such as reinforcement learning (Hoeller et al., 2021; Wijmans et al., 2019; Lee et al., 2024; Ye et al., 2021) and imitation learning (Sadat et al., 2020; Shah et al., 2023b; Loquercio et al., 2021). Despite these advancements, significant challenges persist. Reinforcement learning-based methods often suffer from sample inefficiency and difficulties in directly processing raw, dense sensor inputs, such as depth images. Without sufficient guidance during training, reinforcement learning struggles to converge on an optimal policy that generalizes well across various scenarios or environments. Conversely, imitation learning relies heavily on the availability and quality of labeled trajectories. Obtaining these labeled trajectories is particularly challenging for robotic systems that operate under diverse dynamics models, thereby limiting their broad applicability in flexible robotic systems.
@@
[…]
@@
During training, we leverage the AutoDiff capabilities provided by PyTorch (Paszke et al., 2019) to solve the BLO in IL. For the LL trajectory optimization, we adopt the cubic spline interpolation implementation provided by PyPose (Wang et al., 2023). It supports differentiable batched closed-form solutions, enabling a faster training process. Additionally, the ESDF cost map is convoluted with a Gaussian kernel, enabling a smoother optimization space. This setup allows the loss to be directly backpropagated to update the network parameters in a single step, rather than requiring iterative adjustments. As a result, the UL network and the trajectory optimization can mutually improve each other, enabling a self-supervised learning process.
@@ 4.2 First-order Optimization
We next illustrate the scenario that the LL cost $L$ in (1b) uses first-order optimizers such as GD. Because GD is a simple differentiable iterative method, one can leverage unrolled optimization listed in Algorithm 1 to solve BLO via AutoDiff. It has been theoretically proved by Ji et al. (2021, 2022a) that when the LL problem is strongly convex and smooth, unrolled optimization with LL gradient descent can approximate the hypergradients $\nabla_{\bm{\theta}}$ and $\nabla_{\bm{\gamma}}$ up to an error that decreases linearly with the number of GD steps. For the implicit differentiation, we could leverage the optimality conditions of the LL optimization to compute the implicit gradient. To be specific, using Chain rule and optimality of ${\bm{\mu}}^{*}$ and ${\bm{\nu}}^{*}$, we have the stationary points
@@
A similar linear system can be derived for computing $\nabla_{\bm{\gamma}}U$. Then, the practical methods should first use a first-order optimizer to solve the lower-level problem in (1b) to obtain approximates $\hat{\bm{\mu}}$ and $\hat{\bm{\nu}}$ of the solutions ${\bm{\mu}}^{*}$ and ${\bm{\nu}}^{*}$, which are then incorporated into (30e) to obtain an approximate ^$\widehat{\nabla}_{\bm{\theta}}U$ of the upper-level gradient $\nabla_{\bm{\theta}}U$ (similarly for ^$\widehat{\nabla}_{\bm{\gamma}}U$). Then, the upper-level gradient approximates ^$\widehat{\nabla}_{\bm{\theta}}U$ and ^$\widehat{\nabla}_{\bm{\gamma}}U$ are used to optimize the target variables ${\bm{\theta}}$ and ${\bm{\gamma}}$.
@@
^$\widehat{\nabla}_{\bm{\theta}}U=\frac{\partial U}{\partial f}\frac{\partial f}{\partial{\bm{\theta}}},\quad\widehat{\nabla}_{\bm{\gamma}}U=\frac{\partial U}{\partial M}\frac{\partial M}{\partial{\bm{\gamma}}},$ (32)
@@
To achieve theoretical guarantees, one sufficient condition is that the LL function $L$ in (1b) is twice differentiable and strongly convex respect to $\phi$. The UL function $U$ can be possibly non-convex. Based on (Ji et al., 2021, Theorem 1), the proposed first-order method, which solves the linear system in (31) and uses $\hat{\bm{\mu}}$ and $\hat{\bm{\nu}}$ of the solutions ${\bm{\mu}}^{*}$ and ${\bm{\nu}}^{*}$, can achieve the first-order $\epsilon$-accurate stationary optimality, i.e., $\|\nabla_{\psi}\|\leq\epsilon$, if we use $\log\frac{1}{\epsilon}$-level steps for the approximation in (32). Moreover, when the UL function $U$ in (19) is strongly convex with respect to $\theta$ and $\gamma$, we can further show that the obtained solution achieves the global optimality, based on (Ji and Liang, 2021, Theorem 9).
@@
Given few known facts, logical reasoning aims at deducting the truth value of unknown facts with formal logical rules (Iwańska, 1993; Overton, 2013; Halpern, 2013). Different types of logic have been invented to address distinct problems from various domains, including propositional logic (Wang et al., 2019), linear temporal logic (Xie et al., 2021), and first-order-logic (FOL) (Cropper and Morel, 2021). Among them, FOL decomposes a logic term into lifted predicates and variables with quantifiers, which has strong generalization power and can be applied to arbitrary entities with any compositions. Due to the strong capability of generalization, FOL has been widely leveraged in knowledge graph reasoning (Yang and Song, 2019) and robotic task planning (Chitnis et al., 2022). However, traditional FOL requires a human expert to carefully design the predicates and rules, which is tedious. Automatically summarizing the FOL predicates and rules is a long-standing problem, which is known as inductive logic programming (ILP). However, existing works study ILP only in simple structured data like knowledge graphs. To extend this research into robotics, we will explore how IL can make ILP work with high-dimensional data like RGB images.
@@
One stream of the solutions to ILP is based on forward search algorithms (Cropper and Morel, 2021; Shindo et al., 2023; Hocquette et al., 2024). For example, Popper constructs answer set programs based on failures, which can significantly reduce the hypothesis space (Cropper and Morel, 2021). However, as FOL is a combinatorial problem, search-based methods can be extremely time-consuming as the samples scale up. Recently, some works introduced neural networks to assist the search process (Yang et al., 2017; Yang and Song, 2019; Yang et al., 2022b) or directly implicitly represent the rule (Dong et al., 2019). To name a few, NeurlLP re-formulates the FOL rule inference into the multi-hop reasoning process, which can be represented as a series of matrix multiplications (Yang et al., 2017). Thus, learning the weight matrix becomes equivalent to inducting the rules. Neural logic machines, on the other hand, designed a new FOL-inspired network structure, where the rules are implicitly stored in the network weights (Dong et al., 2019).
Despite their promising results in structured symbolic datasets, such as binary vector representations in BlocksWorld (Dong et al., 2019) and knowledge graphs (Yang et al., 2017), their capability of handling high-dimensional data like RGB images is rarely explored.
@@
To address this gap, we verify the IL framework with the visual action prediction (VAP) task in the LogiCity (Li et al., 2024), a recent logical reasoning benchmark. In VAP, a model must simultaneously discover traffic rules, identify the concept of agents and their spatial relationships, and predict the agents’ actions. As shown in Figure 4, we utilize a grounding network $f$ to predict the agent concepts and spatial relationships, which takes the observations such as images as the model input. The predicted agent concepts and their relationships are then sent to the reasoning module for rule induction and action prediction. The learned rules from the reasoning process are stored in the memory and retrieved when necessary. For example, the grounding network may output concepts “IsTiro($A_{1}$)”, “IsPedestrian($A_{2}$),” and their relationship “NextTo($A_{1}$, $A_{2}$)”. Then the reasoning engine applies the learned rules, “Slow($X$$\leftarrow$($\exists$ $Y$ IsTiro($X$) $\land$ IsPedestrian($Y$) $\land$ NextTo($X$, $Y$)) $\lor$ …”, stored in the memory module to infer the next actions “Slow($A_{1}$)” or “Normal($A_{2}$)”, as is displayed in Figure 5. Finally, the predicted actions and the observed actions from the grounding networks are used for loss calculation. In summary, we formulate this pipeline in (33) and refer to it as imperative logical reasoning (iLogic):
@@
Specifically, in the experiments we use the same cross-entropy function for the UL and LL costs, i.e., $U\doteq L\doteq-\sum a_{i}^{o}\log a_{i}^{p}$, where $a_{i}^{o}$ and $a_{i}^{p}$ are the observed and predicted actions of the $i$-th agent, respectively. This results in a self-supervised simultaneous grounding and rule induction pipeline for logical reasoning. Additionally, $f$ can be any grounding networks and we use a feature pyramid network (FPN) (Lin et al., 2017) as the visual encoder and two MLPs for relationships; $g$ can be any logical reasoning engines and we use a neural logical machine (NLM) (Dong et al., 2019) with parameter ${\bm{\mu}}$; and $M$ can be any memory modules storing the learned rules and we use the MLPs in the NLM parameterized by ${\bm{\nu}}$. More details about the task, model, and list of concepts are presented in the Section 5.2.

### v7

@@ 4 Applications and Examples
Additionally, since IL is a self-supervised learning framework consisting of three primary components, we have three kinds of learning types. This includes (A) given known (pre-trained or human-defined) reasoning engines such as logical reasoning, physical principles, and geometrical analytics, robots can learn a logic-, physics-, or geometry-aware neural perception system, respectively, in a self-supervised manner; (B) given neural perception systems such as a vision foundation model (kirillov2023segment), robots can discover the world rules, e.g., traffic rules, and then apply the rules to future events; and (C) given a memory system, (e.g., experience, world rules, or maps), robots can simultaneously update the neural system and reasoning engine so that they can adapt to novel environments with a new set of rules. Note that if one of the modules has a good initialization, all the three modules can be further finetuned simultaneously. The five examples will cover all three learning types.
@@ 4.1 Closed-form Solution
where the LL solutions ${\bm{\mu}}^{*}({\bm{\theta}},{\bm{\gamma}})$ and ${\bm{\nu}}^{*}({\bm{\theta}},{\bm{\gamma}})$, that contain the implicit components $\frac{\partial{\bm{\mu}}^{*}}{\partial{\bm{\theta}}}$, $\frac{\partial{\bm{\mu}}^{*}}{\partial{\bm{\gamma}}}$, $\frac{\partial{\bm{\nu}}^{*}}{\partial{\bm{\theta}}}$, and $\frac{\partial{\bm{\nu}}^{*}}{\partial{\bm{\gamma}}}$, can be calculated directly due to the closed-form of $\bm{\mu}$$^{*}$ and $\bm{\nu}$$^{*}$. As a result, standard gradient-based algorithms can be applied in updating $\bm{\theta}$} and $\bm{\gamma}$} with gradients given by (2). In this case, there is no approximation error induced by the LL minimization, in contrast to existing widely-used implicit and unrolled differentiation methods that require ensuring a sufficiently small or decreasing LL optimization error for guaranteeing the convergence (ji2021bilevel).
@@
Many symbolic costs can be effectively addressed through closed-form solutions. For example, both linear quadratic regulator (LQR) (shaiju2008formulas) and Dijkstra’s algorithm (dijkstra1959note) can be solved with a determined optimal solution. To demonstrate the effectiveness of IL for closed-form solutions, we next present two examples in path planning to utilize the neural model for reducing the search and sampling space of symbolic optimization.
@@
The A algorithm is a graph search technique that aims to find the global shortest feasible path between two nodes (hart1968formal). Specifically, A selects the path passing through the next node $n\in G$ that minimizes
@@
where $s(n)$ is the cost from the start node to $n$ and $h(n)$ is a heuristic function that predicts the cheapest path cost from $n$ to the goal. The heuristic cost can take the form of various metrics such as the Euclidean and Chebyshev distances. It can be proved that if the heuristic function $h(n)$ is admissible and monotone, i.e., $\forall m\in G,h(n)\leq h(m)+d(m,n)$, where $d(m,n)$ is the cost from $m$ to $n$, A is guaranteed to find the optimal path without searching any node more than once (dechter1985generalized). Due to this optimality, A became one of the most widely used methods in path planning (paden2016surveymotionplanning; smith2012dual; algfoor2015comprehensive).
However, A encounters significant limitations, particularly in its requirement to explore a large number of potential nodes. This exhaustive search process can be excessively time-consuming, especially for low-power robot systems. To address this, recent advancements showed significant potential for enhancing efficiency by predicting a more accurate heuristic cost map using data-driven methods (choudhury2018data; yonetani2021path; kirilenko2023transpath). Nevertheless, these algorithms utilize optimal paths as training labels, which face challenges in generalization, leading to a bias towards the patterns observed in the training datasets.
Figure 2: The framework of iA* search. The network predicts a confined search space, leading to overall improved efficiency. The A search algorithm eliminates the label dependence, resulting in a self-supervised path planning framework.
@@
To address this limitation, we leverage the IL framework to remove dependence on path labeling. Specifically, we utilize a neural network $f$ to estimate the $h$ value of each node, which can be used to reduce the search space. Subsequently, we integrate a differentiable A module, serving as the symbolic reasoning engine $g$ in (1b), to determine the most efficient path. This results in an effective framework depicted in Figure 2, which we refer to as imperative A (iA*) algorithm. Notably, the iA* framework can operate on a self-supervised basis inherent in the IL framework, eliminating the need for annotated labels.
Specifically, the iA* algorithm is to minimize
@@
As shown in (2), once the gradient calculation is completed, we backpropagate the cost $U$ directly to the network $f$. This process is facilitated by the closed-form solution of the A algorithm. For the sake of simplicity, we employ the differentiable A algorithm as introduced by yonetani2021path. It transforms the node selection into an argsoftmax operation and reinterprets node expansion as a series of convolutions, leveraging efficient implementations and AutoDiff in PyTorch (paszke2019pytorch).
@@
End-to-end local planning, which integrates perception and planning within a single model, has recently attracted considerable interest, particularly for its potential to enable efficient inference through data-driven methods such as reinforcement learning (hoeller2021learning; wijmans2019dd; lee2024learning; ye2021auxiliary) and imitation learning (sadat2020perceive; shah2023gnm; loquercio2021learning). Despite these advancements, significant challenges persist. Reinforcement learning-based methods often suffer from sample inefficiency and difficulties in directly processing raw, dense sensor inputs, such as depth images. Without sufficient guidance during training, reinforcement learning struggles to converge on an optimal policy that generalizes well across various scenarios or environments. Conversely, imitation learning relies heavily on the availability and quality of labeled trajectories. Obtaining these labeled trajectories is particularly challenging for robotic systems that operate under diverse dynamics models, thereby limiting their broad applicability in flexible robotic systems.
@@
where $D(\bm{\mu},\bm{z})$ measures the path continuity based on the 1st-order derivative of the path and $A(\bm{u})$ calculates the path smoothness based on 2nd-order derivative. Specifically, we employ the cubic spline interpolation which has a closed-form solution to ensure this continuity and smoothness. This is also essential for generating feasible and efficient paths. On the other hand, the UL cost $U$ is defined as:
@@
During training, we leverage the AutoDiff capabilities provided by PyTorch (paszke2019pytorch) to solve the BLO in IL. For the LL trajectory optimization, we adopt the cubic spline interpolation implementation provided by PyPose (wang2023pypose). It supports differentiable batched closed-form solutions, enabling a faster training process. Additionally, the ESDF cost map is convoluted with a Gaussian kernel, enabling a smoother optimization space. This setup allows the loss to be directly backpropagated to update the network parameters in a single step, rather than requiring iterative adjustments. As a result, the UL network and the trajectory optimization can mutually improve each other, enabling a self-supervised learning process.
@@ 4.2 First-order Optimization
We next illustrate the scenario that the LL cost $L$ in (1b) uses first-order optimizers such as GD. Because GD is a simple differentiable iterative method, one can leverage unrolled optimization listed in Algorithm 1 to solve BLO via AutoDiff. It has been theoretically proved by ji2021bilevel; ji2022will that when the LL problem is strongly convex and smooth, unrolled optimization with LL gradient descent can approximate the hypergradients $\nabla_{\bm{\theta}}$ and $\nabla_{\bm{\gamma}}$ up to an error that decreases linearly with the number of GD steps. For the implicit differentiation, we could leverage the optimality conditions of the LL optimization to compute the implicit gradient. To be specific, using Chain rule and optimality of $\bm{\mu}$$^{*}$ and $\bm{\nu}$$^{*}$, we have the stationary points
@@
A similar linear system can be derived for computing $\nabla_{\bm{\gamma}}U$. Then, the practical methods should first use a first-order optimizer to solve the lower-level problem in (1b) to obtain approximates $\hat{\bm{\mu}}$ and $\hat{\bm{\nu}}$ of the solutions $\bm{\mu}$$^{*}$ and $\bm{\nu}$$^{*}$, which are then incorporated into (31e) to obtain an approximate $\widehat{\nabla}_{\bm{\theta}}U$ of the upper-level gradient $\nabla_{\bm{\theta}}U$ (similarly for $\widehat{\nabla}_{\bm{\gamma}}U$). Then, the upper-level gradient approximates $\widehat{\nabla}_{\bm{\theta}}U$ and $\widehat{\nabla}_{\bm{\gamma}}U$ are used to optimize the target variables $\bm{\theta}$} and $\bm{\gamma}$}.
@@
$\widehat{\nabla}_{\bm{\theta}}U=\frac{\partial U}{\partial f}\frac{\partial f}{\partial{\bm{\theta}}},\quad\widehat{\nabla}_{\bm{\gamma}}U=\frac{\partial U}{\partial M}\frac{\partial M}{\partial{\bm{\gamma}}},$ (33)
@@
To achieve theoretical guarantees, one sufficient condition is that the LL function $L$ in (1b) is twice differentiable and strongly convex respect to $\phi$. The UL function $U$ can be possibly non-convex. Based on (ji2021bilevel, Theorem 1), the proposed first-order method, which solves the linear system in (32) and uses $\hat{\bm{\mu}}$ and $\hat{\bm{\nu}}$ of the solutions $\bm{\mu}$$^{*}$ and $\bm{\nu}$$^{*}$, can achieve the first-order $\epsilon$-accurate stationary optimality, i.e., $\|\nabla_{\psi}\|\leq\epsilon$, if we use $\log\frac{1}{\epsilon}$-level steps for the approximation in (33). Moreover, when the UL function $U$ in (20) is strongly convex with respect to $\theta$ and $\gamma$, we can further show that the obtained solution achieves the global optimality, based on (ji2021lower, Theorem 9).
@@
Given few known facts, logical reasoning aims at deducting the truth value of unknown facts with formal logical rules (iwanska1993logical; overton2013competence; halpern2013thought). Different types of logic have been invented to address distinct problems from various domains, including propositional logic (wang2019satnet), linear temporal logic (xie2021embedding), and first-order-logic (FOL) (cropper2021popper). Among them, FOL decomposes a logic term into lifted predicates and variables with quantifiers, which has strong generalization power and can be applied to arbitrary entities with any compositions. Due to the strong capability of generalization, FOL has been widely leveraged in knowledge graph reasoning (yang2019nlil) and robotic task planning (chitnis2022nsrt). However, traditional FOL requires a human expert to carefully design the predicates and rules, which is tedious. Automatically summarizing the FOL predicates and rules is a long-standing problem, which is known as inductive logic programming (ILP). However, existing works study ILP only in simple structured data like knowledge graphs. To extend this research into robotics, we will explore how IL can make ILP work with high-dimensional data like RGB images.
@@
One stream of the solutions to ILP is based on forward search algorithms (cropper2021popper; shindo2023aILP; cropper2024maxsynth). For example, Popper constructs answer set programs based on failures, which can significantly reduce the hypothesis space (cropper2021popper). However, as FOL is a combinatorial problem, search-based methods can be extremely time-consuming as the samples scale up. Recently, some works introduced neural networks to assist the search process (yang2017neurallp; yang2019nlil; yang2022logicdef) or directly implicitly represent the rule (dong2019nlm). To name a few, NeurlLP re-formulates the FOL rule inference into the multi-hop reasoning process, which can be represented as a series of matrix multiplications (yang2017neurallp). Thus, learning the weight matrix becomes equivalent to inducting the rules. Neural logic machines, on the other hand, designed a new FOL-inspired network structure, where the rules are implicitly stored in the network weights (dong2019nlm).
Despite their promising results in structured symbolic datasets, such as binary vector representations in BlocksWorld (dong2019nlm) and knowledge graphs (yang2017neurallp), their capability of handling high-dimensional data like RGB images is rarely explored.
@@
[…]
@@
Specifically, in the experiments we use the same cross-entropy function for the UL and LL costs, i.e., $U\doteq L\doteq-\sum a_{i}^{o}\log a_{i}^{p}$, where $a_{i}^{o}$ and $a_{i}^{p}$ are the observed and predicted actions of the $i$-th agent, respectively. This results in a self-supervised simultaneous grounding and rule induction pipeline for logical reasoning. Additionally, $f$ can be any grounding networks and we use a feature pyramid network (FPN) (lin2017fpn) as the visual encoder and two MLPs for relationships; $g$ can be any logical reasoning engines and we use a neural logical machine (NLM) (dong2019nlm) with parameter $\bm{\mu}$}; and $M$ can be any memory modules storing the learned rules and we use the MLPs in the NLM parameterized by $\bm{\nu}$}. More details about the task, model, and list of concepts are presented in the Section 5.2.